A small business employs 98 people. All of the employees work for 8 hours a day, 5 days a week. Eleven of the employees receive a wage of $45 an hour and the rest receive $30 an hour. In total, how much does the business pay its employees each week?

Answer by JulietG(1812)

That must be an amazingly profitable business to pay so well!
Each employee works 40 hours(8*5)
11 of those employees make 40*$45, or $1800/week
11 * 1800 = 19,800
The other 87 employees (98-11=87) make 40*$30, or $1200/week
87 * 1200 = 104,400
Total: 19,800 + 104,400 = $124,200